OpenStudy (anonymous):

What is the lateral area of a rectangular prism if the base edges are 6 meters and 4 meters and the height is 7 meters?

OpenStudy (anonymous):

lateral area of a rectangular prism L=perimeter*Height L=(2*b1+2b2)H L=(2*6 + 2*4)*7=____?

OpenStudy (anonymous):

i got it now thanx:)

OpenStudy (anonymous):

YW have fun and good luck .... :D

OpenStudy (anonymous):

thanx